# Grounding Fallacies Misrepresenting Scientific Publications in Evidence

### v1

@@ Abstract
Health-related misinformation claims often falsely cite a credible biomedical publication as evidence, which superficially appears to support the false claim. The publication does not really support the claim, but a reader could believe it thanks to the use of logical fallacies. Here, we aim to detect and to highlight such fallacies, which requires carefully assessing the exact content of the misrepresented publications. To achieve this, we introduce MissciPlus, an extension of the fallacy detection dataset Missci. MissciPlus builds on Missci by grounding the applied fallacies in real-world passages from misrepresented studies. This creates a realistic test-bed for detecting and verbalizing these fallacies under real-world input conditions, and enables novel passage-retrieval tasks. MissciPlus is the first logical fallacy dataset which pairs the real-world misrepresented evidence with incorrect claims, identical to the input to evidence-based fact-checking models. With MissciPlus, we i) benchmark retrieval models in identifying passages that support claims only when fallacies are applied, ii) evaluate how well LLMs articulate fallacious reasoning from misrepresented scientific passages, and iii) assess the effectiveness of fact-checking models in refuting claims that misrepresent biomedical research. Our findings show that current fact-checking models struggle to use relevant passages from misrepresented publications to refute misinformation. Moreover, these passages can mislead LLMs into accepting false claims as true. [^1] https://github.com/UKPLab/arxiv2024-missciplus
@@
Max Glockner, Yufang Hou♢♣, Preslav Nakovand Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), TU Darmstadt and Hessian Center for AI (hessian.AI) IBM Research, Ireland, MBZUAI www.ukp.tu-darmstadt.de
@@ 1 Introduction
Health-related misinformation has caused significant harm in our society (Zarocostas, 2020). Human fact-checking (HFC), which is time-consuming, must prioritize the most impactful claims and struggles with the rapid spread of misinformation (Arnold, 2020; Vosoughi et al., 2018). Two main approaches can automatically combat health-related misinformation. Scientific automated fact-checking (AFC) retrieves scientific evidence documents to support or to refute a claim, but faces challenges with regard to mismatched specificity (Wadden et al., 2022), and the reliance on the availability of refuting evidence (Glockner et al., 2022). Reason-checking (Musi et al., 2023) rejects claims that use fallacious reasoning, which is particularly relevant as often facts are misrepresented or skewed (Brennen et al., 2020). Most logical fallacy detection datasets require that the fallacious reasoning is explicitly stated in the text (Jin et al., 2022; Alhindi et al., 2022), and cannot detect implicit distortions from misused evidence.
@@
The Missci dataset (Glockner et al., 2024a) addresses implicit fallacious reasoning by reconstructing the fallacious logical argument, explicitly verbalizing the flawed reasoning that led to the incorrect claim. Figure 1 shows (parts of) a fallacious logical argument from Missci for the claim that “hydroxychloroquine is a cure for COVID-19”. Each claim in Missci has a kernel of truth which is anchored in the study’s content (“Chloroquine reduced infection of the coronavirus”), denoted as the accurate premise in green. However, additional information from the study undermines the claim’s validity and reveals a reasoning gap (“The study used cell cultures for their experiments”), denoted as publication context in orange. To bridge the reasoning gap between the study content and the claim, the model must verbalize a second premise (“The results can be transferred to humans because the human body consists of cell structure.”), denoted as fallacious premise in blue. The model must also predict the type of fallacy associated with this fallacious premise.
The accurate premise and publication context in Missci constitute manually paraphrased information based on the HFC articles, which are unavailable for new claims in the wild. In a real-world scenario, models must first (i) identify the relevant passages within the entire misrepresented study and (ii) verbalize the fallacious reasoning based on the original scientific text. To address this, we present MissciPlus4), an extension of Missci with 2,257 human-annotated links between the paraphrased information from Missci and the real-world passages from the publications (see the right and the bottom of Figure 1). MissciPlus is the first fallacy dataset to pair real-world misinformation claims with publication passages as evidence. This setup, identical to the input used by AFC models, allows for the evaluation of evidence-based scientific AFC models. Since the claimant based their claim on these evidence passages, they serve as critical evidence in the fact-checking process (Silverman, 2016).
We leverage MissciPlus to answer the following research questions:
@@
    How well can existing ranking approaches select the correct evidence passages (§5)?
@@
    How well can LLMs reconstruct fallacious arguments over original studies compared to paraphrased information in Missci6)?
@@
    Can evidence-based scientific AFC models use the content of the misrepresented publications to detect the claims as misinformation (§7)?
Our findings suggest that lexical/semantic similarity based ranking models perform the best at identifying correct evidence passages. However, all models cannot leverage these passages to refute misinformation.
In summary, we contribute a new dataset MissciPlus to ground fallacies in real-world evidence, which bridges the gap to AFC. We propose novel task definitions, along with extensive experiments on retrieving the required passages of the misrepresented study, and experiments using AFC models and LLMs in detecting science distortions.
@@ 2 Related Work
research focused on surface-level fallacies (Habernal et al., 2017, 2018; Da San Martino et al., 2019; Sahai et al., 2021; Piskorski et al., 2023; Salman et al., 2023). Others extended these inventories to include logical fallacies that may require additional context for detection. All that research relied on educational examples, fake news websites (Jin et al., 2022), or fact-checking articles (Musi et al., 2022; Alhindi et al., 2022) and assumed that the explicitly stated text was sufficient to detect the fallacies. Missci (Glockner et al., 2024a) developed models to verbalize the implicit fallacious reasoning. MissciPlus differs from existing fallacy datasets as it grounds the implicit fallacies in the original content of the real-world evidence documents.
Automated scientific fact-checking
(Wadden et al., 2020; Saakyan et al., 2021; Sarrouti et al., 2021; Kotonya and Toni, 2020; Lu et al., 2023; Vladika and Matthes, 2023, 2024) used scientific documents as evidence to assess the veracity of claims. These approaches face challenges with fine-grained differences, such as specificity mismatches (Wadden et al., 2022). Our work bridges the gap between scientific AFC and fallacy detection and sheds light on the abilities of AFC models to reason over claims and distorted evidence passages.
@@
research (Augenstein, 2021) compares claims and the cited evidence document across various dimensions such as claim strength (Li et al., 2017), certainty of the used language (Pei and Jurgens, 2021), sentence-level causal exaggerations (Yu et al., 2020), quantification of the information match (Wright et al., 2022), or combinations of multiple dimensions (Wührl et al., 2024). Our work differs due to the focus on harmful misinformation and tasks that retrieve the passages and verbalize the fallacious reasoning.
@@ 3 Background
Missci (Glockner et al., 2024a) comprises 208 health-related claims that misrepresent biomedical publications. Inspired by Cook et al. (2018), each claim is modeled as a fallacious logical argument where a claim is the conclusion of its premises. Formally, each argument comprises exactly one accurate premise ($p_{0}$) based on which the claim was made, and at least one fallacious premise $\overline{p}_{i}$ with the fallacy class $f_{i}$ (cf. §A for details of all nine fallacy classes), that is needed to (falsely) conclude the claim $\overline{c}$ from the study’s content. We refer to the pair of fallacious premise and the applied fallacy class ($\overline{p}_{i}$, $f_{i}$) as fallacy. Each fallacy in Missci is linked to one publication context ($s_{i}$) that points to a reasoning gap in the argument, and necessitates the fallacy. Each publication context (and accurate premise $p_{0}$) faithfully summarizes parts of the study. Missci defines the argument reconstruction task as: For each publication context ($s_{i}$) of the argument, given $s_{i}$, the incorrect claim ($\overline{c}$) and the accurate premise ($p_{0}$), the model must verbalize the fallacious premise and detect the applied fallacy class ($\overline{p}_{i}$, $f_{i}$). Due to overlapping fallacy definitions (Bonial et al., 2022), sometimes different interchangeable fallacies ($\overline{p}_{i}$, $f_{i}$) can bridge the same reasoning gap (expressed via $s_{i}$), and only one of them is needed to reconstruct the argument; we refer to Glockner et al. (2024a) for more detail.
@@
In order to investigate the antiviral properties of chloroquine on SARS-CoV after the initiation of infection, Vero E6 cells were infected with the virus and fresh medium supplemented with various concentrations of chloroquine was added immediately after virus adsorption. Infected cells were incubated for an additional 16-18 h, after which the presence of virus antigens was analyzed by indirect immunofluorescence analysis. When chloroquine was added after the initiation of infection, there was a dramatic dose-dependant decrease in the number of virus antigen-positive cells (Fig. 2A). As little as 0.1-1 muM chloroquine reduced the infection by 50% and up to 90-94% inhibition was observed with 33-100 muM concentrations (Fig. 2B). At concentrations of chloroquine in excess of 1 muM, only a small number of individual cells were initially infected, and the spread of the infection to adjacent cells was all but eliminated. A half-maximal inhibitory effect was estimated to occur at 4.4 ± 1.0 muM chloroquine (Fig. 2C). These data clearly show that addition of chloroquine can effectively reduce the establishment of infection and spread of SARS-CoV if the drug is added immediately following virus adsorption. (Vincent et al., 2005).
@@
Figure 2: A real-world passage communicates the paraphrased content $s_{1}$ from Missci that the study used cell cultures for their experiments.
Linked Passages
@@
In Missci, the accurate premise ($p_{0}$) and publication contexts ($s_{i}$) were directly paraphrased from HFC articles and not from the original misrepresented study. Since these HFC articles were specifically written to explain to non-experts why the misrepresented study does not support the claim, the paraphrased publication contexts often reveal the reasoning gaps easily. This severely limits the applicability of Missci in the real world, where models must first identify relevant content from the misrepresented study and reason through complex scientific text. To address this, we created MissciPlus, which links the paraphrased information from Missci with actual passages from the misrepresented study. For example, Figure 2 shows a verbatim passage from the misrepresented study in MissciPlus, which is linked (i.e., communicates the same content) to the paraphrased content that “the study used cell cultures for their experiments” from Missci (see Figure 2). Formally, given a misrepresented study $S$ with its passages $S_{j}\in S$, we linked the passage $S_{j}$ to the corresponding paraphrased information ($p_{0}$ and $s_{i}$) if (parts of) $S_{j}$ entail the paraphrased information ($p_{0}$ or $s_{i}$). Any passage $S_{j}$ linked to $p_{0}$ (or $s_{i}$) can replace $p_{0}$ (or $s_{i}$) during the argument reconstruction. The same passage may be linked to multiple paraphrased information and vice versa. We denote a passage $S_{j}$ as $S_{j}^{0}$ if it links to $p_{0}$.
Subtasks
@@
We consider three sub-tasks for reconstructing fallacious arguments in the wild. First (§5.1), the model must retrieve a passage $S_{j}^{0}$, upon which the claim is based. This is crucial for understanding the general reasoning of the claim. Second (§5.2), the model must retrieve all additional passages $S_{j}$ required to detect fallacies, i.e., passages linked to any publication context $s_{i}$. Lastly (§6), the argument reconstruction task is adapted from Missci, but replaces the paraphrased content with the respective linked passages. We always assume oracle input to establish a strong foundation for each sub-task, laying the groundwork for a more robust end-to-end system in the future.
4 Grounding Missci with Evidence
Data Collection
We selected all fallacious logical arguments from Missci, for which the misrepresented study is openly available via PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/. This resulted in 118 arguments misrepresenting 100 distinct publications, which we automatically split into the constituent passages (cf. §B.1). We used the IMS model from Wright et al. (2022) to pre-select relevant passages and avoid exhaustively annotating every paraphrased information with every passage. The IMS model quantifies the information match between statements and publications, which aligns well with our needs during dataset creation. For each paraphrased information ($p_{0}$ and $s_{i}$), we selected the top-ranked passage according to IMS (cf. §B.2). We collected a minimum of six passages per argument, if possible.
Annotation
@@ 4 Grounding Missci with Evidence
We employed two biology master’s students with annotation experience in biomedical misinformation. The annotators assigned an entailment label to all combinations of selected passage ($S_{j}$) and paraphrased information ($p_{0}$ and $s_{i}$), determining whether the passage $S_{j}$ entails (and hence is linked to) the paraphrased information. Following Glockner et al. (2024b), the annotators could express uncertainty if the entailment relation was ambiguous. If the paraphrased information was not linked to any passage, a corresponding passage was manually selected from the entire study and added to the passages for annotation. After removing four arguments for which no paraphrased information could be linked to any passage, the annotation resulted in 2,257 double-annotated entailment relations between passages and paraphrased information across 114 arguments. We retained the same instances for validation (30 arguments) and test (84 arguments) splits as in Missci. The inter-annotator agreement, measured by Cohen’s $\kappa$, was 0.602. For details about the annotation process please refer to §B.3. Overall, 400 pieces of paraphrased information (88.6% of the accurate premises $p_{0}$; 72.0% of the publication contexts $s_{i}$; 76.8% overall) could be linked to at least one corresponding passage (analysed in §B.4). We give provide descriptive statistics about the passages in §C.

### v2

@@ Abstract
Health-related misinformation claims often falsely cite a credible biomedical publication as evidence. These publications only superficially seem to support the false claim, when logical fallacies are applied. In this work, we aim to detect and to highlight such fallacies, which requires assessing the exact content of the misrepresented publications. To achieve this, we introduce MissciPlus, an extension of the fallacy detection dataset Missci. MissciPlus extends Missci by grounding the applied fallacies in real-world passages from misrepresented studies. This creates a realistic test-bed for detecting and verbalizing fallacies under real-world input conditions, and enables new and realistic passage-retrieval tasks. MissciPlus is the first logical fallacy dataset which pairs the real-world misrepresented evidence with incorrect claims, identical to the input to evidence-based fact-checking models. With MissciPlus, we i) benchmark retrieval models in identifying passages that support claims only with fallacious reasoning, ii) evaluate how well LLMs verbalize fallacious reasoning based on misrepresented scientific passages, and iii) assess the effectiveness of fact-checking models in refuting claims that misrepresent biomedical research. Our findings show that current fact-checking models struggle to use misrepresented scientific passages to refute misinformation. Moreover, these passages can mislead LLMs into accepting false claims as true. [^1] Code and data are available at https://github.com/UKPLab/naacl2025-missciplus
@@
Max Glockner, Yufang Hou♢♡, Preslav Nakovand Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), TU Darmstadt and Hessian Center for AI (hessian.AI) IT:U Interdisciplinary Transformation University Austria IBM Research Ireland, MBZUAI www.ukp.tu-darmstadt.de
@@ 1 Introduction
Health-related misinformation has caused significant harm in our society (Zarocostas, 2020). Human fact-checking (HFC), which is time-consuming, must prioritize the most impactful claims and struggles with the rapid spread of misinformation (Arnold, 2020; Vosoughi et al., 2018). Two main approaches can automatically combat health-related misinformation: Scientific automated fact-checking (AFC) retrieves scientific evidence documents to support or to refute a claim, but faces challenges with mismatched specificity between claim and evidence (Wadden et al., 2022), and the reliance on the availability of refuting evidence (Glockner et al., 2022). Reason-checking (Musi et al., 2023) rejects claims that use fallacious reasoning, which is particularly relevant since facts are often misrepresented or skewed (Brennen et al., 2020). Unlike most logical fallacy detection datasets, which assume that fallacious reasoning is explicitly stated (Jin et al., 2022; Alhindi et al., 2022), real-world scenarios often involve fallacies that are not explicitly articulated.
The Missci dataset (Glockner et al., 2024a) addresses implicit fallacious reasoning by reconstructing the fallacious logical argument, explicitly verbalizing the fallacies that led to the incorrect claim. However, Missci provides pre-selected and simplified short phrases (as shown in Figure 1) instead of the actual misrepresented study as evidence. In a real-world scenario the relevant parts of the misrepresented study are not presented in this form. Here, the models must first (i) identify the relevant passages within the entire misrepresented study and (ii) verbalize the fallacious reasoning based on the original scientific text. To address this, we present MissciPlus4), an extension of Missci with 2,257 human-annotated links between the simplified phrases from Missci and the real-world passages from the misrepresented publications.
@@
Figure 1 shows (parts of) a fallacious logical argument from Missci for the claim that “hydroxychloroquine is a cure for COVID-19”. Each claim in Missci has a kernel of truth which is anchored in the study’s content (“Chloroquine reduced infection of the coronavirus”), denoted as the accurate premise in green. However, additional information from the same study (“The study used cell cultures for their experiments”) undermines the claim’s validity and reveals a reasoning gap. The relevant content from the study that is needed to identify this reasoning gap between claim and study is referred to as the publication context. The reasoning gap between the study’s content and the claim as a conclusion indicates the fallacious reasoning. To bridge this reasoning gap the model must verbalize a fallacious premise (“The results can be transferred to humans because the human body consists of cell structure.”), blue in Figure 1, and classify the fallacy associated with this fallacious premise (“fallacy of composition”). Unlike Missci, where the accurate premise and publication context are provided directly, MissciPlus requires LLMs to retrieve the required real-world passages from the misrepresented study and reason over them. A complete fallacious argument, including the misrepresented study passages, is provided in §B. MissciPlus is the first fallacy dataset to pair real-world misinformation claims with publication passages as evidence. This setup is identical to the input used by evidence-based scientific AFC models and enables their evaluation over real-world fallacious misinformation. We use MissciPlus to answer the following research questions:
@@
    How well can existing ranking approaches select the required publication passages (§5)?
@@
    How well can LLMs reconstruct fallacious arguments using original studies compared to the simplified content in Missci6)?
@@
    Can evidence-based scientific AFC models use the content of misrepresented publications to detect the claims as misinformation (§7)?
Our findings suggest that lexical and semantic similarity based ranking models perform the best at identifying the required evidence passages. However, all models cannot leverage these passages to refute misinformation. In summary, we contribute MissciPlus, an extension of Missci to ground fallacies in real-world evidence, which bridges the gap to AFC. We propose novel task definitions, along with extensive experiments on retrieving the required passages from the misrepresented study needed to detect misrepresented publications in the wild, as well as experiments using AFC models and LLMs in detecting science distortions.
@@ 2 Related Work
Much research on fallacy detection has primarily focused on surface-level fallacies (Habernal et al., 2017, 2018; Da San Martino et al., 2019; Sahai et al., 2021; Piskorski et al., 2023; Salman et al., 2023). Other works extended these inventories to include logical fallacies that may require additional context for detection. However, all of this research relied on educational examples, fake news websites (Jin et al., 2022), or fact-checking articles (Musi et al., 2022; Alhindi et al., 2022) and assumed that the explicitly stated text was sufficient to detect the fallacies. Missci (Glockner et al., 2024a) developed models to verbalize the implicit fallacious reasoning. MissciPlus differs from existing fallacy datasets by grounding implicit fallacies in real-world evidence documents.
Scientific AFC
A large body of research on scientific AFC used scientific documents as evidence to assess the veracity of claims (Wadden et al., 2020; Saakyan et al., 2021; Sarrouti et al., 2021; Kotonya and Toni, 2020; Lu et al., 2023; Vladika and Matthes, 2023, 2024). These approaches face challenges with fine-grained differences, such as specificity mismatches (Wadden et al., 2022). Our work bridges the gap between scientific AFC and fallacy detection and sheds light on the abilities of AFC models to reason over claims and misrepresented evidence passages.
@@
A very related research area concerns science communication research (Augenstein, 2021), which compares claims and the cited evidence document across various dimensions such as claim strength (Li et al., 2017), certainty of the used language (Pei and Jurgens, 2021), sentence-level causal exaggerations (Yu et al., 2020), quantification of the information match (Wright et al., 2022), or combinations of multiple dimensions (Wuehrl et al., 2024). Our work differs by focusing on harmful misinformation and tasks that involve retrieving relevant passages and articulating the fallacious reasoning.
@@ 3.1 Preliminaries
Missci (Glockner et al., 2024a) comprises 208 inaccurate health-related claims that misrepresent scientific publications. Each claim is modeled as a fallacious logical argument where a claim is the (wrong) conclusion of its premises. Formally, each argument comprises exactly one accurate premise ($p_{0}$) based on which the claim was made, and at least one fallacious premise $\overline{p}_{i}$ with the fallacy class $f_{i}$ (cf. §A for details of all nine fallacy classes), that is needed to (falsely) conclude the claim $\overline{c}$ from the study’s content. We refer to the pair of fallacious premise and the applied fallacy class ($\overline{p}_{i}$, $f_{i}$) as fallacy. Each fallacy in Missci is linked to one publication context ($s_{i}$). The publication context contains the content of the misrepresented study, that necessitates the fallacy. For example, knowing that the study’s observations were limited to cell cultures (publication context, orange, Figure 1) reveals that the claim about the effectiveness (in humans) is unjustified. We define this disconnect between study content and its purported conclusions as a reasoning gap, which fallacious premises attempt to bridge. Each publication context (and accurate premise $p_{0}$) faithfully summarizes parts of the study. Missci defines the argument reconstruction task as: Given the publication context ($s_{i}$), the incorrect claim ($\overline{c}$) and the accurate premise ($p_{0}$), the model must verbalize the fallacious premise and detect the applied fallacy class ($\overline{p}_{i}$, $f_{i}$).
@@
In order to investigate the antiviral properties of chloroquine on SARS-CoV after the initiation of infection, Vero E6 cells were infected with the virus and fresh medium supplemented with various concentrations of chloroquine was added immediately after virus adsorption. Infected cells were incubated for an additional 16-18 h, after which the presence of virus antigens was analyzed by indirect immunofluorescence analysis. When chloroquine was added after the initiation of infection, there was a dramatic dose-dependant decrease in the number of virus antigen-positive cells (Fig. 2A). As little as 0.1-1 muM chloroquine reduced the infection by 50% and up to 90-94% inhibition was observed with 33-100 muM concentrations (Fig. 2B). At concentrations of chloroquine in excess of 1 muM, only a small number of individual cells were initially infected, and the spread of the infection to adjacent cells was all but eliminated. A half-maximal inhibitory effect was estimated to occur at 4.4 ± 1.0 muM chloroquine (Fig. 2C). These data clearly show that addition of chloroquine can effectively reduce the establishment of infection and spread of SARS-CoV if the drug is added immediately following virus adsorption.
@@
Figure 2: A real-world passage (Vincent et al., 2005) communicates the paraphrased content $s_{1}$ from Missci that the study used cell cultures for their experiments.
3.2 Linked Passages
@@
In Missci, the accurate premise ($p_{0}$) and publication contexts ($s_{i}$) were manually paraphrased from the HFC article and not the study itself. Since these HFC articles were specifically written to explain to non-experts why the misrepresented study does not support the claim, the paraphrased publication contexts often reveal the reasoning gaps easily. This severely limits the applicability of Missci in the real world, where models must identify relevant content from the entire misrepresented study, and reason through complex scientific text. To address this, MissciPlus links the simplified paraphrased information from Missci with actual passages from the misrepresented study. Figure 2 shows a verbatim passage from the misrepresented study in MissciPlus, which is linked (i.e., communicates the same content) to the paraphrased content that “the study used cell cultures for their experiments” from Missci (see Figure 1). Formally, given a misrepresented study $S$ with its passages $S_{j}\in S$, we linked the passage $S_{j}$ to the corresponding paraphrased information ($p_{0}$ and $s_{i}$) if (parts of) $S_{j}$ entail the paraphrased information ($p_{0}$ or $s_{i}$). Any passage $S_{j}$ linked to $p_{0}$ (or $s_{i}$) can replace $p_{0}$ (or $s_{i}$) during the argument reconstruction. The same passage may be linked to multiple paraphrased information and vice versa. We denote a passage $S_{j}$ as $S_{j}^{0}$ if it links to $p_{0}$.
3.3 Subtasks
@@
We consider three sub-tasks for reconstructing fallacious arguments in the wild. First (§5.1), the model must retrieve a passage $S_{j}^{0}$, upon which the claim is based. This is crucial for understanding the general reasoning of the claim. Second (§5.2), the model must retrieve all additional passages $S_{j}$ required to detect fallacies, i.e., passages linked to any publication context $s_{i}$. Lastly (§6), the argument reconstruction task is adapted from Missci, but replaces the paraphrased content with the respective linked passages. In reality, each subtask relies on the output of preceding subtasks. In this work, we aim to establish a strong foundation for each sub-task individually and assume oracle input for each, laying the groundwork for a more robust end-to-end system in the future.
4 Grounding MissciPlus with Evidence
@@
To create MissciPlus, we selected all fallacious logical arguments from Missci, for which the full misrepresented study is available via PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/. This resulted in 118 fallacious arguments misrepresenting 100 distinct publications, which we automatically split into the constituent passages (cf. §C.1). We used the IMS model (Wright et al., 2022) to pre-select relevant passages and avoid exhaustively annotating every paraphrased information with every passage. The IMS model quantifies the information match between textual statements and scientific text, which aligns well with our needs. For each paraphrased information ($p_{0}$ and $s_{i}$), we selected the top-ranked passage according to IMS (cf. §C.2) and collected a minimum of six passages per argument in total, if possible. We employed two biology master’s students with annotation experience in biomedical misinformation on Missci. The annotators assigned an entailment label by comparing each paraphrased information ($p_{0}$ and $s_{i}$) with each selected passage ($S_{j}$), determining whether $S_{j}$ entails (and hence is linked to) the paraphrased information. Following Glockner et al. (2024b), the annotators could express uncertainty if the entailment relation was ambiguous. If the paraphrased information was not linked to any pre-selected passage after consolidation, one annotator manually selected a corresponding passage from the entire study, if possible, which was then double-annotated. We removed four arguments for which no paraphrased information could be linked to any passage, yielding 2,257 double-annotated relations between passages and paraphrased information across 114 arguments. We retained the same instances for validation (30 arguments) and test (84 arguments) splits as in Missci. The inter-annotator agreement, measured by Cohen’s $\kappa$, was 0.602. For details about the annotation process please refer to §C.3. Overall, 400 pieces of paraphrased information (88.6% of the accurate premises $p_{0}$; 72.0% of the publication contexts $s_{i}$; 76.8% overall) could be linked to at least one passage (analysed in §C.4). Descriptive statistics about the passages are in §D.